please help me if someone can please
IF(px+q)^2=9x^2+kx+16,what is the value of k?OPTIONS ARE
(A)3
(B)4
(C)12
(D)24
(E)30

Answer by ewatrrr(24785)

(px+q)^2=9x^2+kx+16
p^2x^2 + 2pqx + q^2 = 9x^2+kx+16
p^2 = 9 , p = ±3
q^2 = 16 , q = ± 4
...
2pq = k
(D) k = 24 works